# A Deep Learning Method for Simultaneous Denoising and Missing Wedge Reconstruction in Cryogenic Electron Tomography

### v2

@@ Abstract
Cryogenic electron tomography is a technique for imaging biological samples in 3D. A microscope collects a series of 2D projections of the sample, and the goal is to reconstruct the 3D density of the sample called the tomogram. Reconstruction is difficult as the 2D projections are noisy and can not be recorded from all directions, resulting in a missing wedge of information. Tomograms conventionally reconstructed with filtered back-projection suffer from noise and strong artifacts due to the missing wedge. Here, we propose a deep-learning approach for simultaneous denoising and missing wedge reconstruction called DeepDeWedge. The algorithm requires no ground truth data and is based on fitting a neural network to the 2D projections using a self-supervised loss. DeepDeWedge performs better than CryoCARE and IsoNet, which are state-of-the-art methods for denoising and missing wedge reconstruction, and similarly and, in some cases, better than the combination of the two methods. At the same time, DeepDeWedge is simpler than this two-step approach, as it does denoising and missing wedge reconstruction simultaneously rather than sequentially.
@@ 1 Introduction
Cryogenic electron tomography (cryo-ET) is a powerful cryo-electron microscopy (cryo-EM) technique for obtaining 3D models of biological samples such as cells, viruses, and proteins. An important application of cryo-ET is visualizing biological macromolecules like proteins in situ, i.e., in their (close-to) native environment. This preserves biological context, which can greatly improve the understanding of the tasks and workings of the macromolecules [WB16].
In cryo-ET, the sample to be imaged is first prepared on a grid and then frozen. Next, a transmission electron microscope records a tilt series, which is a collection of 2D projections of the sample’s 3D scattering potential or density. Each projection in the tilt series is recorded after tilting the sample for a fixed number of degrees around the microscope’s tilt axis.
From this tilt series, a tomogram, i.e., a discretized estimate of the sample’s 3D density, can be estimated using computational techniques. For this inverse problem, numerous approaches have been proposed. The most commonly used tomographic reconstruction technique is filtered back-projection (FBP) [Rad88, WB16, TB20].
Two major obstacles limit the resolution and interpretability of tomograms reconstructed with FBP and similar methods:
@@
    A missing wedge of information: The range of angles at which useful images can be collected is often limited to, for example, ±$\pm 60^{\circ}$ rather than the desired full range of ±$\pm 90^{\circ}$. This is due to the increased thickness of the sample in the direction of the electron beam for tilt angles of large magnitude [Liu+22]. The missing data in the tilt series is wedge-shaped in the Fourier domain.
@@
    Noisy data: The total electron dose during tilt series acquisition must be low because biological samples studied with cryo-ET are usually sensitive to radiation damage. Thus, the individual projections of the tilt series have low contrast and a low signal-to-noise ratio (SNR).
In this paper, we propose DeepDeWedge, a deep learning-based imaging approach that simultaneously performs denoising and missing wedge reconstruction. The algorithm takes a single tilt series as input and aims to estimate a noise-free 3D reconstruction of the samples’ density with a filled-in missing wedge.
The algorithm can also be applied to a small dataset of tilt-series, typically two to ten, that were collected from samples of the same type, for example, instances of the same cell or solutions containing purified particles, in which case it estimates the noise-free, missing-wedge-corrected 3D densities of all the samples.
DeepDeWedge fits a randomly initialized network with a self-supervised loss for simultaneous denoising and missing wedge reconstruction. After fitting, the network is applied to the same data to estimate the tomogram. DeepDeWedge only uses the tilt series of the density we wish to reconstruct and no other training data.
Related methods are effective for denoising and missing wedge reconstruction individually but not simultaneously. Specifically, IsoNet [Liu+22], a very related deep learning approach for denoising and missing wedge reconstruction, does well at missing wedge reconstruction, but its denoising performance is low compared to state-of-the-art denoising methods [Mal+23]. The current state-of-the-art approaches for denoising in cryo-ET are inspired by noise2noise [Leh+18], a framework for deep-learning-based image denoising. Popular software packages that implement noise2noise-based denoising methods for cryo-ET tomograms are CryoCARE [Buc+19], Topaz [Bep+20] and Warp [TC19]. DeepDeWedge is most related to such noise2noise-based denoising approaches and to IsoNet. We discuss the exact relations in Section 2.1.2 after describing our algorithm.
DeepDeWedge performs on par with IsoNet on a pure missing wedge reconstruction problem and achieves state-of-the-art denoising performance on a pure denoising problem. Moreover, we find that the performance of DeepDeWedge for the typical joint denoising and wedge reconstruction problem is equal and, in some cases, better than the two-step approach of applying IsoNet to tomograms denoised with a state-of-the-art noise2noise-like denoiser. Moreover, DeepDeWedge is simpler and requires fewer hyperparameters to tune than the two-step approach.
Our method is also conceptually related to un-trained neural networks, which reconstruct an image or volume based on fitting a neural network to data [UVL18, HH18]. Un-trained networks also only rely on fitting a neural network to given measurements. However, they rely on the bias of convolutional neural networks towards natural images [HS20a, HS20], whereas in our setup, we fit a network on measurement data to be able to reconstruct from the same measurements.
For cryo-EM-related problems other than tomographic reconstruction, deep learning approaches for missing data reconstruction and denoising have also recently been proposed. [Zha+23, ] proposed a method to restore the state of individual particles inside tomograms, and [LHZ23] proposed a variant of IsoNet to resolve the preferred orientation problem in single-particle cryo-EM [LHZ23].
Another line of works for cryo-ET reconstruction considers domain-specific tomographic reconstruction methods that incorporate prior knowledge of biological samples into the reconstruction process to compensate for missing wedge artifacts, for example, ICON [Den+16], and MBIR [Yan+19]. For an overview of such reconstruction methods, we refer readers to the introductory sections of works by [Din+19, ] and [BBC22, ]. [Liu+22, ] found that IsoNet outperforms both ICON and MBIR.
The rest of this paper is structured as follows: In the remainder of this section, we provide background on cryo-ET. In Section 2.1, we describe the DeepDeWedge algorithm, and Section 2.2 contains experimental results on real and synthetic cryo-ET data.
Tilt Series Acquisition Model and Problem Statement
@@
A tilt series $\mathbf{t}=\big{(}\mathbf{t}_{-K},\ldots,\mathbf{t}_{K}\big{)}$ is a collection of 2D projections, where each 2D projection $\mathbf{t}_{k}$ is a measurement of an underlying ground truth volume $\mathbf{v}^{*}$ obtained with the electron microscope. The relation of the 2D projections and the ground truth volume is as follows: By the Fourier slice theorem (see e.g. [Mal93]), the 2D Fourier transform $\mathbf{F}\mathbf{t}_{k}$ of each projection image $\mathbf{t}_{k}$ of a tilt series is a noisy observation of a 2D central slice through true volume’s 3D Fourier transform $\mathbf{F}\mathbf{v}^{*}$ multiplied with an additional filter, i.e.,
@@
In Equation (1), the rotation operator $\mathbf{R}_{k}$ spatially rotates the 3D Fourier transform $\mathbf{F}\mathbf{v}^{*}$ of the volume $\mathbf{v}^{*}$ by the tilt angle $\alpha_{k}$ around the microscope’s tilt axis. Then, the slice-operator $\mathbf{S}$ extracts the central 2D slice of the volume’s Fourier transform that is perpendicular to the microscope’s optical axis. The filter $\mathbf{C}_{k}$ is the contrast transfer function (CTF) of the microscope and models optical aberrations. The term $\mathbf{F}\mathbf{\mathbf{n}}_{k}$ is the Fourier transform of a random 2D image-domain noise term $\mathbf{\mathbf{n}}_{k}$. It is often assumed that the noise $\mathbf{\mathbf{n}}_{k}$ comes from a Poisson distribution (shot noise). Another common assumption is that the Fourier-domain noise $\mathbf{F}\mathbf{\mathbf{n}}_{k}$ is Gaussian, but not necessarily white [BBS20, Fra21]. In this work, we make the weaker assumption that the 2D noise terms $\mathbf{\mathbf{n}}_{k}$ and $\mathbf{\mathbf{n}}_{\ell}$ of any two distinct projections indexed with $k$ and $\ell$ ($k\neq\ell$) have zero means and are independent.
@@
Equation (1) is commonly used as a Fourier-domain model of the image formation process in cryo-EM [Sig16, BBS20]. As the Fourier slice theorem assumes continuous representations of volumes and images, this model is exact only in the continuous case. In this work, we consider discrete representations of volumes and images, as is common in cryo-EM practice. Thus, Equation (1) holds only approximately.
@@
The range of tilt angles $\alpha_{k}$ that yield useful projections $\mathbf{t}_{k}$ is typically limited to, e.g., ±$\pm 60^{\circ}$ rather than the full range ±$\pm 90^{\circ}$. Therefore, Equation (1) implies that there is a wedge-shaped region of the sample’s Fourier representation $\mathbf{F}\mathbf{v}^{*}$ which is not covered by any of the Fourier slices $\mathbf{F}\mathbf{t}_{k}$.
@@
We consider the problem of reconstructing a tomogram, i.e., a discretized estimate of a sample’s density from a noisy, incomplete tilt series. This is a difficult inverse problem due to the missing wedge and the high noise level.
@@
2 Results
We now describe our algorithm for estimating missing-wedge corrected, denoised cryo-ET tomograms and evaluate its effectiveness on real and simulated data.
2.1 Denoising and missing wedge reconstruction algorithm (DeepDeWedge)
@@
Step 1: Data preparation. First, split the tilt series $\mathbf{t}$ into two sub-tilt-series ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}}$ with the even/odd split or the frame-based split. The even/odd split partitions the tilt series into even and odd projections based on their order of acquisition. The frame-based split can be applied if the tilt series is collected using dose fractionation and entails averaging only the even and odd frames recorded at each tilt angle. We recommend the frame-based splitting approach whenever possible. After splitting, we have two sub-tilt-series.
Next, reconstruct both sub-tilt-series independently with FBP and apply CTF correction. This yields a pair of two coarse reconstructions $\big{(}\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}}),\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})\big{)}$ of the sample’s 3D density. Finally, use a 3D-sliding window procedure to extract $N$ overlapping cubic sub-tomogram pairs $\{({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{1}}})\}_{i=1}^{N}$from both FBP reconstructions. The size and number $N$ of these sub-tomogram cubes is a hyperparameter. Experiments on synthetic data presented in Appendix E.1 suggest that larger sub-tomograms tend to yield better results, up to a point.
Step 2: Model fitting. Fit a randomly initialized network $f_{\text{$\boldsymbol{\theta}$}}$, we use a U-Net [RFB15], with weights $\boldsymbol{\theta}$ by repeating the following steps until convergence:
@@
  • Update the model: Update the model weights $\boldsymbol{\theta}$ by performing gradient steps to minimize the per-sample loss
@@
    $\ell\left(\text{$\boldsymbol{\theta}$},i\right)=\left|\left|\left(\mathbf{M}\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}+2\mathbf{M}^{C}\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}\right)\mathbf{F}\left(f_{\text{$\boldsymbol{\theta}$}}\big{(}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}\big{)}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}.$ (2)
@@
    Here, $\mathbf{M}_{\text{$\boldsymbol{\varphi}$}_{i}}$ is the rotated version of the wedge mask $\mathbf{M}$ and $\mathbf{M}^{C}:=\mathbf{I}-\mathbf{M}$ is the complement of the mask. For the gradient updates, we use the Adam optimizer [KB15] and perform a single pass through the $N$ model input and target sub-tomograms generated before.
@@
  • Update the missing wedges of the model inputs: For each $i=1,\ldots,N$, update the missing wedge the i-th sub-tomogram ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}$ produced in Step 1 by passing it through the current model $f_{\text{$\boldsymbol{\theta}$}}$, and inserting the predicted content of the missing wedge, as follows
@@
    ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\leftarrow\mathbf{F}^{-1}\Big{(}\mathbf{M}\mathbf{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}+\mathbf{M}^{C}\mathbf{F}f_{\text{$\boldsymbol{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\right)\Big{)}.$ (3)
@@
    In the next input-target generation step, the model inputs $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}\}_{i=1}^{N}$ are constructed using the updated sub-tomograms. We do not update the missing wedges of the sub-tomograms $\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{1}}}\}_{i=1}^{N}$used to generate the model targets, since their missing wedges are masked out in the per-sample loss, c.f. Equation (2), and therefore play no role in model fitting.

### v3

@@ Abstract
Cryogenic electron tomography is a technique for imaging biological samples in 3D. A microscope collects a series of 2D projections of the sample, and the goal is to reconstruct the 3D density of the sample called the tomogram. Reconstruction is difficult as the 2D projections are noisy and can not be recorded from all directions, resulting in a missing wedge of information. Tomograms conventionally reconstructed with filtered back-projection suffer from noise and strong artifacts due to the missing wedge. Here, we propose a deep-learning approach for simultaneous denoising and missing wedge reconstruction called DeepDeWedge. The algorithm requires no ground truth data and is based on fitting a neural network to the 2D projections using a self-supervised loss. DeepDeWedge is simpler than current state-of-the-art approaches for denoising and missing wedge reconstruction, performs competitively and produces more denoised tomograms with higher overall contrast.
@@ 1 Introduction
Cryogenic electron tomography (cryo-ET) is a powerful cryo-electron microscopy (cryo-EM) technique for obtaining 3D models of biological samples such as cells, viruses, and proteins. An important application of cryo-ET is visualizing biological macromolecules like proteins in situ, i.e., in their (close-to) native environment. Imaging in situ preserves biological context, which can greatly improve the understanding of the workings of macromolecules [1].
In cryo-ET, the sample to be imaged is first prepared on a grid and then frozen. Next, a transmission electron microscope records a tilt series, which is a collection of 2D projections of the sample’s 3D scattering potential or density. Each projection in the tilt series is recorded after tilting the sample for a number of degrees around the microscope’s tilt axis.
From this tilt series, a tomogram, i.e., a discretized estimate of the sample’s 3D density, can be estimated using computational techniques. For this inverse problem, numerous approaches have been proposed. The most commonly used tomographic reconstruction technique is filtered back-projection (FBP) [2, 1, 3]. Two major obstacles limit the resolution and interpretability of tomograms reconstructed with FBP and similar methods:
@@
    Noisy data: The total electron dose during tilt series acquisition must be low because biological samples are sensitive to radiation damage. Thus, the individual projections of the tilt series have low contrast and a low signal-to-noise ratio (SNR).
@@
    A missing wedge of information: The range of angles at which useful images can be collected is often limited to, for example, ±$\pm 60^{\circ}$ rather than the desired full range of ±$\pm 90^{\circ}$. This is due to the increased thickness of the sample in the direction of the electron beam for tilt angles of large magnitude [4]. The missing data in the tilt series is wedge-shaped in the Fourier domain.
@@
To define the missing wedge and motivation for our work, we now state a model for the tilt series acquisition process in cryo-ET: A tilt series $\mathbf{t}=\big{(}\mathbf{t}_{-K},\ldots,\mathbf{t}_{K}\big{)}$ is a collection of 2D projections, where each 2D projection $\mathbf{t}_{k}$ is a measurement of an underlying ground truth volume $\mathbf{v}^{*}$ obtained with the electron microscope. By the Fourier slice theorem (see e.g. [5]), the 2D Fourier transform $\mathbf{F}\mathbf{t}_{k}$ of each projection image $\mathbf{t}_{k}$ of a tilt series is a noisy observation of a 2D central slice through true volume’s 3D Fourier transform $\mathbf{F}\mathbf{v}^{*}$ multiplied with an additional filter, i.e.,
@@
In Equation (1), the rotation operator $\mathbf{R}_{k}$ spatially rotates the 3D Fourier transform $\mathbf{F}\mathbf{v}^{*}$ of the volume $\mathbf{v}^{*}$ by the tilt angle $\alpha_{k}$ around the microscope’s tilt axis. Then, the slice-operator $\mathbf{S}$ extracts the central 2D slice of the volume’s Fourier transform that is perpendicular to the microscope’s optical axis. The filter $\mathbf{C}_{k}$ is the contrast transfer function (CTF) of the microscope and models optical aberrations. The term $\mathbf{F}\mathbf{\mathbf{n}}_{k}$ is the Fourier transform of a random 2D image-domain noise term $\mathbf{\mathbf{n}}_{k}$. It is often assumed that the noise $\mathbf{\mathbf{n}}_{k}$ comes from a Poisson distribution (shot noise). Another common assumption is that the Fourier-domain noise $\mathbf{F}\mathbf{\mathbf{n}}_{k}$ is Gaussian, but not necessarily white [6, 7]. In this work, we make the weaker assumption that the 2D noise terms $\mathbf{\mathbf{n}}_{k}$ and $\mathbf{\mathbf{n}}_{\ell}$ of any two distinct projections indexed with $k$ and $\ell$ ($k\neq\ell$) have zero means and are independent.
@@
Equation (1) is commonly used as a Fourier-domain model of the image formation process in cryo-EM [8, 6]. As the Fourier slice theorem assumes continuous representations of volumes and images, this model is exact only in the continuous case. In this work, we consider discrete representations of volumes and images, as is common in cryo-EM practice. Thus, Equation (1) holds only approximately.
@@
The range of tilt angles $\alpha_{k}$ that yield useful projections $\mathbf{t}_{k}$ is typically limited to, e.g., ±$\pm 60^{\circ}$ rather than the full range ±$\pm 90^{\circ}$. Therefore, by Equation (1) there is a wedge-shaped region of the sample’s Fourier representation $\mathbf{F}\mathbf{v}^{*}$ which is not covered by any of the Fourier slices $\mathbf{F}\mathbf{t}_{k}$.
@@
We consider the problem of reconstructing a tomogram, i.e., a discretized estimate of a sample’s density from a noisy, incomplete tilt series. This is a challenging inverse problem due to the high noise level and the missing wedge.
Recently, deep-learning-based methods for denoising and missing wedge reconstruction have been proposed. However, these are effective for denoising and missing wedge reconstruction individually but not simultaneously. Specifically, IsoNet [4], a closely related work, does well at missing wedge reconstruction, but its denoising performance is low compared to state-of-the-art denoising methods [9]. The current state-of-the-art approaches for denoising in cryo-ET build on Noise2Noise [10], a framework for deep-learning-based image denoising. Popular software packages that implement Noise2Noise-based denoising methods for cryo-ET tomograms are CryoCARE [11], Topaz [12] and Warp [13].
In this paper, we propose DeepDeWedge, a deep learning-based approach for tomogram reconstruction that simultaneously performs well on denoising and missing wedge reconstruction.
DeepDeWedge takes one (or more; see Section 2) tilt series as input and aims to estimate a noise-free 3D reconstruction of the samples’ density with a filled-in missing wedge. To achieve this, we propose fitting a randomly initialized network with a self-supervised loss for simultaneous denoising and missing wedge reconstruction. After fitting, the network is applied to the same data to estimate the tomogram. DeepDeWedge only uses the tilt series of the density we wish to reconstruct and no other training data. DeepDeWedge is most related to Noise2Noise-based denoising approaches and to IsoNet, both introduced above. We discuss the exact relations in Section 2.2 after describing our algorithm. DeepDeWedge performs on par with IsoNet on a pure missing wedge reconstruction problem and achieves state-of-the-art denoising performance on a pure denoising problem. Moreover, we find that the performance of DeepDeWedge for the joint denoising and wedge reconstruction problem is similar to the two-step approach of applying IsoNet to tomograms denoised with a state-of-the-art Noise2Noise-like denoiser. Moreover, DeepDeWedge is simpler and requires fewer hyperparameters to tune than the two-step approach.
@@
2 DeepDeWedge: Denoising and Missing Wedge Reconstruction Algorithm
@@
  1. 1.
    Data preparation: First, split the tilt series $\mathbf{t}$ into two sub-tilt-series ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}}$ with the even/odd split or the frame-based split. The even/odd split partitions the tilt series into even and odd projections based on their order of acquisition. The frame-based split can be applied if the tilt series is collected using dose fractionation and entails averaging only the even and odd frames recorded at each tilt angle. We recommend the frame-based splitting approach whenever possible. After splitting, we have two sub-tilt-series.
    Next, reconstruct both sub-tilt-series independently with FBP and apply CTF correction. This yields a pair of two coarse reconstructions $\big{(}\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}}),\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})\big{)}$ of the sample’s 3D density. from both FBP reconstructions. The size and number $N$ of these sub-tomogram cubes is a hyperparameter. Experiments on synthetic data presented in the supplementary information suggest that larger sub-tomograms tend to yield better results up to a point.
  2. 2.
    Model fitting: Fit a randomly initialized network $\mathrm{f}_{\text{$\bm{\theta}$}}$, we use a U-Net [14], with weights $\bm{\theta}$ by repeating the following steps until convergence:
    1. (a)
@@
      Update the model: Update the model weights $\bm{\theta}$ by performing gradient steps to minimize the per-sample loss
@@
      $\ell\left(\text{$\bm{\theta}$},i\right)=\left|\left|\left(\mathbf{M}\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}+2\mathbf{M}^{C}\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}\right)\mathbf{F}\left(\mathrm{f}_{\text{$\bm{\theta}$}}\big{(}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}\big{)}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{1}}}\right)\right|\right|_{2}^{2}.$ (2)
@@
      Here, $\mathbf{M}_{\text{$\bm{\varphi}$}_{i}}$ is the rotated version of the wedge mask $\mathbf{M}$ and $\mathbf{M}^{C}:=\mathbf{I}-\mathbf{M}$ is the complement of the mask. For the gradient updates, we use the Adam optimizer [15] and perform a single pass through the $N$ model input and target sub-tomograms generated before.
    3. (c)
@@
      Update the missing wedges of the model inputs: For each $i=1,\ldots,N$, update the missing wedge the i-th sub-tomogram ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}$ produced in Step 1 by passing it through the current model $\mathrm{f}_{\text{$\bm{\theta}$}}$, and inserting the predicted content of the missing wedge, as follows
@@
      ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\leftarrow\mathbf{F}^{-1}\Big{(}\mathbf{M}\mathbf{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}+\mathbf{M}^{C}\mathbf{F}\mathrm{f}_{\text{$\bm{\theta}$}}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{0}}}\right)\Big{)}.$ (3)
@@
      In the next input-target generation step, the model inputs $\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}\}_{i=1}^{N}$ are constructed using the updated sub-tomograms. We do not update the missing wedges of the sub-tomograms $\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{v}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{\text{1}}}\}_{i=1}^{N}$ used to generate the model targets, since their missing wedges are masked out in the per-sample loss, c.f. Equation (2), and therefore play no role in model fitting.
  3. 3.